You are here: Foswiki>AdaptiveOptics Web>Logs>Log_20110607 (09 Jun 2011, GBrusa)Edit Attach
Observer(s): L Busoni, A. Puglisi, S. Esposito, E. Pinna (Arcetri), G. Brusa (LBTO)

Telescope Operator:

SW Support:

Telescope Support: J. Little, M. Midkiff

Instrument Support:

We started the activity around 7am by installing the retroreflector, after some software update checks and a few issues we were able to start working with the AO system around 8:30am.

The issues were related to the fact that apparently the SET command did not apply the correct flat. We had to get around by re-applying the flat thru the Preset window of the GUI since at the point the SET command is no longer available from the main GUI (details missing here...).

The software update were: a) the gain optimization is always done before the end of the PresetAO, so that the system is ready for AO with the optimized gain. b) the slope null vector is now a function of the instrument rotator angle

We tested again the H-band image that had the same feature that before, i.e. a 'dip' in the center of the image.

We measured the NCPA (non common path aberrations) for two more instrument rotator angles 113 and 193 in FeII band. It is now clear that the main non common path aberration is an astigmatism of ~ 60nm rms that rotates with the instrument rotator angle.

We verified that the non common path aberration measured at 193 degrees improves the image in J-band as well, i.e. the aberration seems to be independent form the observing band (SR with and without slope null application: 45.4 and 39.7).

The retro-reflector is left installed since the night observations wil only make use of Lucifer (on SX or left side).

-- Main.GBrusa - 09 Jun 2011
Topic revision: r1 - 09 Jun 2011, GBrusa
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback